OpenStudy (anonymous):

Ill Give Medals!! Given f(x) = x + 2 and g(x) = x2 – 4, find the function (g/f)(x).

jimthompson5910 (jim_thompson5910):

hint: Difference of squares a^2 - b^2 = (a-b)(a+b)

OpenStudy (alexandervonhumboldt2):

so you have as @iambatman and @jim_thompson5910 said you have to simplify this thing: (x+2)/((x+2)*(x-2)) let us know what you get

jimthompson5910 (jim_thompson5910):

iambatman, you have g and f mixed up

OpenStudy (anonymous):

Oops!

OpenStudy (anonymous):

\[\frac{ g(x) }{ f(x) }=\frac{ (x^2-4) }{ (x+2) }\] nice catch jim!

OpenStudy (anonymous):

How do i know that excluded value? I got the answer x-2 for the problem. Im just missing the excluded value..Help Pleasee

OpenStudy (anonymous):

I actually just figured out the excluded value is -2(: Thanks guys!

jimthompson5910 (jim_thompson5910):

you cannot divide by zero so if the denominator is zero, then x+2 = 0 leads to x = -2 like you got